In a circle, the length of an arc with a measure 120 is 14pi. Find the radius of the cirlce
Answer by checkley75(3666)

SET UP A RATIO THE ARC (CIR./ARC) AND THE THE MEASUREMENT (2PIR/14PI).
360/120=2PIR/14PI
3=2R/14
2R=3*14
2R=42
R=42/2
R=21 ANSWER.